OpenStudy (anonymous):

If y varies directly as x, and y = 8 when x = 6, find x when y = 16. A) 2 B) 3 C) 12 D) 14

OpenStudy (anonymous):

8/6 = 16/x x = 16*6/8 = 12

OpenStudy (swissgirl):

hahahah u beat me to it lol

OpenStudy (anonymous):

lol thanks for explaining it to me :) and thx swissgirl lol im loving the name ;D

OpenStudy (swissgirl):

lol